# Shadow in the Cache: Unveiling and Mitigating Privacy Risks of KV-cache in LLM Inference

### v1

@@ Abstract
The Key-Value (KV) cache, which stores intermediate attention computations (Key and Value pairs) to avoid redundant calculations, is a fundamental mechanism for accelerating Large Language Model (LLM) inference. However, this efficiency optimization introduces significant yet underexplored privacy risks. This paper provides the first comprehensive analysis of these vulnerabilities, demonstrating that an attacker can reconstruct sensitive user inputs directly from the KV-cache. We design and implement three distinct attack vectors: a direct Inversion Attack, a more broadly applicable and potent Collision Attack, and a semantic-based Injection Attack. These methods demonstrate the practicality and severity of KV-cache privacy leakage issues. To mitigate this, we propose KV-Cloak, a novel, lightweight, and efficient defense mechanism. KV-Cloak uses a reversible matrix-based obfuscation scheme, combined with operator fusion, to secure the KV-cache. Our extensive experiments show that KV-Cloak effectively thwarts all proposed attacks, reducing reconstruction quality to random noise. Crucially, it achieves this robust security with virtually no degradation in model accuracy and minimal performance overhead, offering a practical solution for trustworthy LLM deployment.
@@ I Introduction
Figure 1: LLM inference service workflow in a privacy-preserving setting and its associated security risks.
However, the storage and potential sharing of the KV-cache introduce significant yet underexplored privacy concerns [38, 36], as illustrated in Figure 1. This vulnerability stems from a critical trade-off made in production systems: while end-user communication with the LLM service is typically encrypted (represented by the black lines in our figure), the KV-cache itself is almost always processed, transmitted between compute nodes, and persisted in plaintext. This design choice is a concession to performance, as the unacceptable latency overhead from cryptographically securing the often gigabyte-scale KV-cache would violate the stringent demands of real-time inference. The feasibility of exploiting such plaintext data exposure is not merely theoretical. It was recently demonstrated by the LeftoverLocals attack [28]. This vulnerability in certain AMD GPUs enables an unprivileged attacker to read data remnants from the GPU’s shared on-chip memory. This provides a concrete method to intercept and exfiltrate plaintext KV-cache data originating from other users’ concurrent inference sessions, proving that sensitive conversational data is indeed at risk. Crucially, the KV-cache is inherently derived from and directly correlated with user inputs. Consequently, if an attacker gains access to the KV-cache, they could potentially reverse-engineer sensitive information in the user’s original prompts or private data [6, 17]. A critical research gap exists in understanding and mitigating the privacy risks associated with KV-caches in LLMs, necessitating further in-depth research.
@@
  • RQ1: Is an attacker able to reconstruct user inputs from the KV-caches?
@@
We first explore two relatively direct attack paths. The first is the KV-cache Inversion Attack, a method that attempts to leverage known model weight matrices to directly reverse-calculate the input from the KV-cache. The second, more general-purpose approach is the KV-cache Collision Attack. This attack reframes input reconstruction as a matching task based on forward computation: an attacker iteratively uses a local model instance to generate KV-caches for candidate inputs and compares them against the intercepted target KV-cache to find a match. Because this method does not rely on any reverse computation, it has broader applicability. Finally, we introduce a novel, semantic-based KV-cache Injection Attack. The core of this attack is to leverage the LLM’s powerful capability to understand and execute instructions. By appending a specific instruction, such as “Repeat the previous content,” to the end of an intercepted KV-cache context, an attacker can induce the model to “echo” or generalize the core information held within the KV-cache.
@@
In response to this challenge, this paper provides an affirmative answer by first analyzing the shortcomings of existing privacy-preserving techniques. Conventional methods like full cryptographic encryption [2] or Homomorphic Encryption (HE) [22] introduce prohibitive computational overhead and latency, rendering them impractical for the high-throughput demands of LLM inference. Meanwhile, applying Differential Privacy (DP) [8] requires adding a level of noise that often severely degrades the model’s inference accuracy to an unacceptable degree. Even specialized solutions like KV-Shield [38], while lightweight, suffer from critical security flaws. Their fixed shuffling mechanism, as we analyzed in Section IV-A, is vulnerable to statistical analysis and is incompatible with modern LLM architectures that use features like Rotary Positional Embedding (RoPE). These limitations highlight the urgent need for a novel defense mechanism, .
Therefore, we propose KV-Cloak, a lightweight and secure mechanism for KV-cache obfuscation designed to overcome these challenges. At its core, KV-Cloak employs a reversible matrix-based obfuscation scheme that guarantees lossless model accuracy. Its security is multi-layered: it applies secret invertible linear transformations to obscure statistical properties, and critically, introduces a one-time random permutation matrix for each data block. This dynamic permutation prevents attackers from building stable algebraic relations across multiple queries. To further enhance performance, KV-Cloak utilizes operator fusion, an optimization where a portion of the secret obfuscation matrices is algebraically fused into the LLM’s attention layer weights offline. This shifts the primary computational cost away from the latency-sensitive online inference phase, striking an effective balance between robust security, lossless accuracy, and high performance.
@@
  • Conducting extensive experiments to evaluate attacks and defenses: We conduct a systematic evaluation to empirically demonstrate and quantify the feasibility of attacks that reconstruct user input from the KV-cache of state-of-the-art LLMs, establishing it as a practical and severe threat. We further prove that our proposed KV-Cloak is a highly practical solution that achieves robust security, near-lossless model fidelity, and high efficiency simultaneously. Our experiments show that KV-Cloak, with its negligible impact on accuracy and minimal latency overhead (mostly $10\%$), successfully resolves the stark trade-off between security and utility that plagues alternative approaches such as DP.
@@ II-A Transformer-based LLM Inference
Prevailing large language models, such as LLaMA [32], DeepSeek [20] and Qwen [37], are predominantly based on the Transformer decoder architecture [42].
Self-Attention Mechanism.
The core of the Transformer architecture is the self-attention mechanism. It captures dependencies between different positions within an input sequence, allowing the model to flexibly focus on various parts of the sequence to better understand the context. For an input sequence $X=(x_{1},\ldots,x_{n})$, the self-attention layer first applies linear transformations to $x_{i}$ at each position to obtain Query ($q$), Key ($k$), and Value ($v$) vectors,
$q_{i}=x_{i}W_{q}^{\top}{R}_{\Theta,i}^{d},\quad k_{i}=x_{i}W_{k}^{\top}{R}_{\Theta,i}^{d},\quad v_{i}=x_{i}W_{v}^{\top}.$ (1)
Here, $W_{q},W_{k},W_{v}$ are learnable weight matrices for the attention layer, and ${R}_{\Theta,i}^{d}$ denotes the Rotary Position Embedding (RoPE)[29]. Then, the self-attention layer computes the dot product of the current query $q_{i}$ with all preceding keys $k_{j}$. The results are scaled, and a softmax function is applied to obtain attention weights $a_{ij}$. These weights are then used to compute a weighted sum of the value vectors. Finally, this sum is passed through a linear projection layer $W_{o}$ to produce the output $o_{i}$ for the current position. This process is detailed in Eq. (2), where $d$ is the dimension of each attention head.
$a_{ij}=\frac{\exp\left(q_{i}{k_{j}}^{\top}/\sqrt{d}\right)}{\sum_{t=1}^{i}\exp\left(q_{i}{k_{t}}^{\top}/\sqrt{d}\right)},o_{i}=\left(\sum_{j=1}^{i}a_{ij}v_{j}\right)W_{o}^{\top}.$ (2)
This mechanism enables the model to dynamically weigh the importance of different parts of the input sequence, thereby capturing complex contextual dependencies.
Autoregressive LLM Inference. The essence of large language model inference is to model the probability distribution $P(x_{1},\ldots,x_{n})$ of a token sequence $x_{1},\ldots,x_{n}$. Due to the sequential nature of natural language, prevailing LLMs predominantly employ an autoregressive text generation approach in inference tasks. The joint probability of the entire sequence can typically be autoregressively decomposed [5] into a product of conditional probabilities,
$P(x_{1},\ldots,x_{n})=P(x_{1})P(x_{2}|x_{1})\cdots P(x_{n}|x_{1},\ldots,x_{n-1}).$ (3)
Specifically, the model generates the output sequence one token at a time. When generating each new token, the model considers not only the initial user-provided input but also all previously generated tokens as context. For example, when generating the token $x_{i}$, the model utilizes information from tokens $x_{1}$ to $x_{i-1}$. While this sequential generation method effectively captures dependencies within the sequence, it introduces significant computational redundancy. At each step of generating a new token, attention computations for all preceding tokens (particularly the calculation of $k$ and $v$) must be recalculated. This repetitive computation substantially impacts inference efficiency.
KV-cache in LLM Inference. In the context of autoregressive generation, the previously generated tokens form the context for predicting the next token. To avoid redundant computations, LLMs employ a caching strategy known as the KV-cache. For each processed token (either from the input or previously generated by the model), its corresponding $k$ and $v$ vectors, computed by the self-attention layers, are stored in the KV-cache. When generating the next token, the model reuses these cached $k$ and $v$ vectors from all preceding tokens, only needing to compute the $q,k,v$ vector for the current prediction step and the $k$ and $v$ vectors for the most recently generated token to add to the cache. This significantly accelerates the inference process, especially for long sequences.
@@ II-B Privacy Attacks against LLMs
Inversion Attacks. While LLMs offer powerful capabilities, their inference process exposes new attack surfaces that can lead to the theft of sensitive user inputs. Privacy attacks targeting the inference process can be broadly categorized into two types: (1) Output-based inversion attacks, which aim to reconstruct inputs from the model’s final outputs (e.g., probability distributions or embedding vectors), and (2) Intermediate state-based inversion attacks, which seek to extract information from the model’s internal computational states [15, 21, 34] (e.g., hidden states). Early work focused on inverting final representations [15], such as sentence embeddings, with methods like Vec2Text [21]. However, these approaches overlooked the richer information in deeper layers. The Embed Parrot attack [34] demonstrated that even deep hidden states can be exploited to accurately reconstruct user inputs, establishing the entire inference pipeline as a critical attack surface.
Unlike a hidden state (the fused output of a Transformer layer), the KV-cache contains the more primitive Key (K) and Value (V) vectors that serve as direct inputs to the attention calculation. Crucially, the KV-cache is a specialized data structure explicitly designed for persistence and reuse across generation steps to optimize inference. This unique role makes it a more potent and exploitable source of privacy leakage than other transient states.
Side-Channel Attacks Exploiting the KV-cache. Currently, research on attacks against the KV-cache has been primarily limited to side-channel attacks. A prominent example is PromptPeek [36], which exploits timing discrepancies created by KV-cache sharing in multi-tenant LLM services. An attacker observes variations in response latency to determine if their guessed prefix has hit another user’s cached entries, allowing them to reconstruct the prompt without direct data access.
However, the viability of such timing side-channel attacks is diminished by modern memory management optimizations. Contemporary inference engines, particularly those using PagedAttention [14], manage the KV-cache in non-contiguous memory blocks (pages), with a typical default block size of 16 tokens. This block-based management means that for a cache hit to occur, an attacker would need to correctly guess an entire block of 16 tokens simultaneously. For an LLM with a vocabulary size in the hundreds of thousands, the search space for a single block exceeds $100,000^{16}$, rendering the attack computationally infeasible. In contrast, our work considers a different, more direct threat model: if an attacker can gain direct access to the KV-cache content (e.g., through memory access or network eavesdropping), they can bypass these side-channel limitations. This direct access enables a more practical collision attack on a per-token basis, reducing the attack complexity to a feasible level and exposing a more fundamental vulnerability. Given that the KV-cache is commonly stored and transmitted in plaintext to preserve performance, this paper aims to systematically explore this underexplored problem of direct, content-based attacks.
III Attack Landscape: Inferring Private User Inputs from KV-cache
While KV-cache is crucial for enhancing the inference efficiency of LLMs, it concurrently introduces novel privacy vulnerabilities. The contextual information stored within KV-cache can contain private user data. Specifically, entries in the KV-cache often correspond directly to tokens in the user’s input sequence. An attacker gaining access to this KV-cache information could potentially infer the original user input, thereby compromising user privacy.
In this section, we aim to study an important research question: is an attacker able to conduct a prompt reconstruction attack to recover the user input from the KV-cache? We first present the threat model of the attack. Subsequently, we propose two distinct attacks to achieve this goal.
@@ III-A Threat Model
In our threat model, we assume that an LLM server has deployed an open-source LLM based service. After a user uses the service with private user inputs, it generates a large number of KV-caches. The attacker gains access to the KV-caches and tries to infer the private inputs, as showed in Figure 2.
Figure 2: The attacker restores user input based on model parameters and leaked KV-cache.
Attacker’s Objective. The primary objective of the attacker is to recover the user input prompt from the accessed KV-cache. In this paper, we assume that the attacker attempts to restore the user input verbatim, which may significantly compromise the privacy of the user.
Attacker’s Capabilities. We consider an attacker with relatively strong yet still practical capabilities, as follows.
  • Access to KV-cache. The attacker can obtain the KV-caches stored by an LLM server. This can be achieved by eavesdropping on the communication channel between nodes (e.g. through the LeftoverLocals attack [28]), reading persisted KV-cache from storage, or other unspecified means of accessing the KV-cache pool.
  • Access to the foundation model on which the deployed LLM is developed. We assume a gray-box setting where the attacker can gain access to the foundation model’s weights on which the deployed LLM is developed [^1] This assumption is realistic given that now most LLM services are developed on open-source foundation models. First, disclosing the underlying foundation model is often a fundamental requirement of open-source licenses. Second, even if model information isn’t publicly disclosed, an attacker could identify the specific foundation model via model fingerprinting attacks [23, 7].. The deployed model could be a fine-tuned version of the foundation model.
@@ III-B Input Reconstruction Attacks from KV-cache
Figure 3: Workflow of input reconstruction attacks.
@@ III-B1 KV-cache Inversion Attack
The KV-cache Inversion Attack, as showed in Figure 3a, is a direct method that attempts to reverse-engineer the input to the attention layer by leveraging the KV-cache data in conjunction with the model’s known attention projection matrices. As defined in the forward computation (Eq. (1)), an attacker who obtains the KV-cache and knows the model’s attention layer parameters $W_{k}$ and $W_{v}$ could theoretically perform a reverse calculation to recover the attention layer’s input $x_{i}$, as Eq. (4).
$x_{i}=k_{i}({R}_{\Theta,i}^{d})^{-1}(W_{k}^{\top})^{-1},\quad x_{i}=v_{i}(W_{v}^{\top})^{-1}.$ (4)
However, the practicality of this direct inversion attack is severely constrained by two fundamental factors. First, it requires the attention projection matrices (e.g., $W_{k},W_{v}$) to be invertible. This condition is met by some earlier models using standard Multi-Head Attention (MHA), such as LLaMA-7B [32], but not by many contemporary models. Architectures like LLaMA 3 [9], Qwen [37], and DeepSeek [20] employ optimizations like Grouped-Query Attention (GQA)[3] or Multi-Head Latent Attention (MLA)[26], which often result in non-square, non-invertible projection matrices, rendering the attack impractical. Second, the attack’s effectiveness is largely confined to the first decoder layer’s KV-cache, as its inversion directly yields the input sequence’s embeddings. In contrast, inverting from deeper layers only recovers intermediate hidden states, from which reconstructing the original text is a significantly more complex and data-intensive problem often requiring auxiliary models [34].
Due to these significant limitations, the KV-cache Inversion Attack, while theoretically straightforward, has a narrow scope of applicability and does not pose a threat to most modern LLM architectures.

### v3

@@ Abstract
The Key-Value (KV) cache, which stores intermediate attention computations (Key and Value pairs) to avoid redundant calculations, is a fundamental mechanism for accelerating Large Language Model (LLM) inference. However, this efficiency optimization introduces significant yet underexplored privacy risks. This paper provides the first comprehensive analysis of these vulnerabilities, demonstrating that an adversary can reconstruct sensitive user inputs directly from the KV-cache. We design and implement three distinct attack vectors: a direct Inversion Attack, a more broadly applicable and potent Collision Attack, and a semantic-based Injection Attack. These methods demonstrate the practicality and severity of KV-cache privacy leakage issues. To mitigate this, we propose KV-Cloak, a novel, lightweight, and efficient defense mechanism. KV-Cloak uses a reversible matrix-based obfuscation scheme, combined with operator fusion, to secure the KV-cache. Our extensive experiments show that KV-Cloak effectively thwarts all proposed attacks, reducing reconstruction quality to random noise. Crucially, it achieves this robust security with virtually no degradation in model accuracy and minimal performance overhead, offering a practical solution for trustworthy LLM deployment.
@@ I Introduction
Figure 1: Overview of the privacy-preserving LLM inference workflow and the associated KV-cache leakage threat model. While user-server communication is encrypted (black lines), the KV-cache is often transmitted and stored in plaintext (red dashed lines), creating a surface for privacy attacks.
However, the storage and potential sharing of the KV-cache introduce significant yet underexplored privacy concerns [41, 39], as illustrated in Figure 1. This vulnerability stems from a critical trade-off made in production systems: while end-user communication with the LLM service is typically encrypted (represented by the black lines), the KV-cache itself is almost always processed, transmitted between compute nodes, and persisted in plaintext. This design choice is a concession to performance, as the unacceptable latency overhead from cryptographically securing the often gigabyte-scale KV-cache would violate the stringent demands of real-time inference. Notably, this risk is amplified in the emerging paradigm of confidential Model-as-a-Service (MaaS). Here, TEE adoption for inference privacy [5, 10] introduces a critical architectural trade-off: high-throughput architectures require deliberately externalizing the massive KV-cache from the TEE’s protection boundary. This architectural design—not a vulnerability—directly exposes the KV-cache to the Cloud Service Provider (CSP), who inherently possesses it without requiring any additional exploit [43].
Crucially, the direct correlation between the KV-cache and user inputs [7, 21], combined with the architectural realities that expose it in plaintext, creates a severe and practical attack vector. A critical research gap therefore exists in mitigating these risks, necessitating a practical and secure defense mechanism.
@@
  • RQ1: Is an adversary able to reconstruct user inputs from the KV-caches?
@@
We first explore two relatively direct attack paths. The first is the KV-cache Inversion Attack, a method that attempts to leverage known model weight matrices to directly reverse-calculate the input from the KV-cache. The second, more general-purpose approach is the KV-cache Collision Attack. This attack reframes input reconstruction as a matching task based on forward computation: an adversary iteratively uses a local model instance to generate KV-caches for candidate inputs and compares them against the intercepted target KV-cache to find a match. Because this method does not rely on any reverse computation, it has broader applicability. Finally, we introduce a novel, semantic-based KV-cache Injection Attack. The core of this attack is to leverage the LLM’s powerful capability to understand and execute instructions. By appending a specific instruction, such as “Repeat the previous content,” to the end of an intercepted KV-cache context, an adversary can induce the model to “echo” or generalize the core information held within the KV-cache.
@@
In response to this challenge, this paper provides an affirmative answer by first analyzing the shortcomings of existing privacy-preserving techniques. Conventional methods like full cryptographic encryption [2] or Homomorphic Encryption (HE) [26] introduce prohibitive computational overhead and latency, rendering them impractical for the high-throughput demands of LLM inference. Meanwhile, applying Differential Privacy (DP) [11] requires adding a level of noise that often severely degrades the model’s inference accuracy to an unacceptable degree. Even specialized solutions like KV-Shield [41], while lightweight, suffer from critical security flaws. Their fixed shuffling mechanism, as we analyzed in Section V-A, is vulnerable to statistical analysis and is incompatible with modern LLM architectures that use features like Rotary Positional Embedding (RoPE)[32]. These limitations highlight the urgent need for a novel defense mechanism.
Therefore, we propose KV-Cloak, a lightweight and secure mechanism for KV-cache obfuscation designed to overcome these challenges. At its core, KV-Cloak employs a reversible matrix-based obfuscation scheme that guarantees lossless model accuracy. Its security is multi-layered: it applies secret invertible linear transformations to obscure statistical properties, and critically, introduces a one-time random permutation matrix for each data block. This dynamic permutation prevents adversaries from building stable algebraic relations across multiple queries. To further enhance performance, KV-Cloak utilizes operator fusion, an optimization where a portion of the secret obfuscation matrices is algebraically fused into the LLM’s attention layer weights offline. This shifts the primary computational cost away from the latency-sensitive online inference phase, striking an effective balance between robust security, lossless accuracy, and high performance.
@@
  • Conducting extensive experiments to evaluate attacks and defenses: We conduct a systematic evaluation to empirically demonstrate and quantify the feasibility of attacks that reconstruct user input from the KV-cache of state-of-the-art LLMs, establishing it as a practical and severe threat. We further prove that our proposed KV-Cloak is a highly practical solution that achieves robust security, near-lossless model fidelity, and high efficiency simultaneously. Our experiments show that KV-Cloak, with its negligible impact on accuracy and minimal latency overhead (mostly $1\%$), successfully resolves the stark trade-off between security and utility that plagues alternative approaches such as DP.
@@ II-A Transformer-based LLM Inference
Prevailing LLMs, such as LLaMA [35], DeepSeek [24] and Qwen [40], are predominantly based on the Transformer decoder architecture [47].
Self-Attention Mechanism. For an input token $x_{i}$, the self-attention layer generates Query ($q_{i}$), Key ($k_{i}$), and Value ($v_{i}$) vectors via linear transformations. Crucially, modern architectures incorporate RoPE, denoted as ${R}_{\Theta,i}^{d}$:
$q_{i}=x_{i}W_{q}^{\top}{R}_{\Theta,i}^{d},\;\;k_{i}=x_{i}W_{k}^{\top}{R}_{\Theta,i}^{d},\;\;v_{i}=x_{i}W_{v}^{\top}.$ (1)
where $W_{(\cdot)}$ are learnable weight matrices. The attention score $a_{ij}$ is computed as the scaled dot product between $q_{i}$ and preceding keys $k_{j\leq i}$. The final output $o_{i}$ is the weighted sum of values projected by $W_{o}$:
$a_{ij}=\frac{\exp(q_{i}{k_{j}}^{\top}/\sqrt{d})}{\sum_{t=1}^{i}\exp(q_{i}{k_{t}}^{\top}/\sqrt{d})},\;\;o_{i}=\left(\textstyle\sum_{j=1}^{i}a_{ij}v_{j}\right)W_{o}^{\top}.$ (2)
Autoregressive LLM Inference. LLM inference models the sequence probability $P(x_{1},\ldots,x_{n})$ through autoregressive decomposition [6]. The generation process predicts tokens sequentially based on the conditional probability:
$P(x_{1},\ldots,x_{n})=\prod_{i=1}^{n}P(x_{i}|x_{<i}).$ (3)
While effective for capturing dependencies, this sequential nature implies that generating $x_{i}$ depends on the full history $x_{<i}$. Without caching, the model must re-compute the Key and Value vectors for all preceding tokens at each step. This computational redundancy creates a substantial bottleneck for real-time inference.
KV-cache in LLM Inference. To mitigate the computational redundancy of autoregressive generation, LLMs utilize the KV-cache mechanism [28]. Instead of recalculating the Key ($K$) and Value ($V$) matrices for the entire context window at each step, the system persists these intermediate states in memory. When generating a new token $x_{i}$, the model only computes the current $q_{i},k_{i},v_{i}$, performs attention between $q_{i}$ and the cached history, and subsequently appends the new $(k_{i},v_{i})$ pair to the cache. This strategy significantly reduces latency but necessitates managing a persistent state proportional to the sequence length.
@@ II-B Privacy Attacks against LLMs
II-B1 Inference-Phase Inversion Attacks
Inference-time privacy attacks are categorized into: (1) Output-based, reconstructing inputs from final probabilities or embeddings [19, 25]; and (2) Intermediate-state-based, exploiting internal representations [37]. While early works like Vec2Text [25] focused on sentence embeddings, Embed Parrot [37] demonstrated that deep hidden states allow for more accurate reconstruction.
However, current research often overlooks the KV-cache. Unlike fused hidden states, the KV-cache retains the primitive Key ($K$) and Value ($V$) vectors—the raw inputs to the attention mechanism. Crucially, to support autoregressive generation, the KV-cache is architecturally designed for persistence, making it a richer and more exploitable surface than transient computational states.
II-B2 Limitations of Side-Channel Attacks
Existing threats to the KV-cache are primarily side-channel attacks, such as PromptPeek [39], which infers prompts by measuring latency variations caused by cache sharing. However, the practicality of such timing attacks is severely mitigated by modern memory management.
Systems like PagedAttention [17] manage the KV-cache in non-contiguous blocks (defaulting to 16 tokens). To trigger a cache hit (and thus a timing signal), an adversary must correctly guess an entire block simultaneously. With a vocabulary size $|V|\approx 10^{5}$, the search space explodes to $|V|^{16}$, rendering brute-force side-channels computationally infeasible. In contrast, we address the direct access threat model. Given that KV-caches are often processed in plaintext for performance, a compromised system allows adversaries to bypass probabilistic guessing and perform precise, per-token collision attacks.
III Attack Landscape: Input Reconstruction from KV-cache
While the KV-cache serves as a cornerstone for efficient LLM inference, it concurrently introduces a critical and underexplored privacy surface. Unlike deep hidden states that represent fused semantic information, entries in the KV-cache maintain a direct, element-wise correspondence to the tokens in the user’s input sequence. This structural characteristic implies that an adversary gaining access to the KV-cache does not merely obtain abstract embeddings, but potentially holds the raw materials to reverse-engineer the original user prompt.
In this section, we investigate the central research question: Is an adversary able to reconstruct user inputs from the KV-cache? We first define a realistic threat model grounded in modern cloud architectures, and subsequently detail three distinct attack vectors designed to exploit this vulnerabilities.
@@ III-A Threat Model
This work focuses on the threat landscape of LLM inference services within a confidential computing paradigm. We consider the inference service provider as the adversary.
Adversary’s Objective. The primary objective of the adversary is to recover the user input prompt from the accessed KV-cache. We target verbatim reconstruction because it represents the most severe breach of privacy, allowing the extraction of exact credentials, PII, or proprietary logic contained in the input.
Adversary’s Capabilities. We assume the adversary can obtain both the KV-cache and the model weights, but does not observe the ephemeral runtime activations within the GPU registers. These assumptions are well-grounded in the realities of current confidential computing inference paradigms:
  • KV-cache Access. The adversary is capable of obtaining the KV-cache stored by the LLM server. This is a realistic assumption, as high-throughput and scalable cloud-native services deliberately externalize the large-scale KV-cache to non-secure memory or persistent storage pools to meet performance demands [43]. We thus consider this access a result of an intentional performance-security trade-off, not a traditional vulnerability.
  • Foundation Model Access. We assume a gray-box setting where the adversary can access the foundation model weights. This access is realistic through two primary paths: (1) For closed-source services, the CSP inherently owns the model and possesses the weights. (2) For services built on open-source models, the provider may be required to disclose the base model for licensing compliance, or the model can be identified using fingerprinting techniques [27, 8].
@@ III-B Input Reconstruction Attacks from KV-cache
Figure 2: Workflow of the three proposed KV-cache input reconstruction attacks.
@@ III-B1 KV-cache Inversion Attack
As illustrated in Figure 2a, the KV-cache Inversion Attack functions as a baseline algebraic adversary. It attempts to mathematically invert the attention mechanism by exploiting the linear projection of Key ($k$) and Value ($v$) vectors.
Attack Formulation. Based on the forward pass definition in Eq. (1), an adversary possessing the plaintext KV-cache and the model weights ($W_{k},W_{v}$) can theoretically recover the input state $x_{i}$ of the attention layer via matrix inversion:
$x_{i}=k_{i}({R}_{\Theta,i}^{d})^{-1}(W_{k}^{\top})^{-1},\;\;x_{i}=v_{i}(W_{v}^{\top})^{-1}.$ (4)
Architectural and Semantic Constraints. While algebraically sound, the feasibility of this attack in production environments is severely limited by two fundamental constraints: (1) The attack strictly requires the projection matrices $W_{k}$ and $W_{v}$ to be square and full-rank invertible. This condition holds for legacy Multi-Head Attention (MHA) [36] architectures (e.g., LLaMA-7B). However, modern State-of-the-Art (SOTA) models, including LLaMA-3 [12], Qwen [40], and DeepSeek [24], adopt efficiency optimizations such as Grouped-Query Attention (GQA) [3] or Multi-Head Latent Attention (MLA) [30]. These mechanisms typically project inputs into lower-dimensional subspaces, resulting in non-square matrices where unique inversion is mathematically impossible (i.e., the problem becomes ill-posed). (2) the attack’s effectiveness is largely confined to the first decoder layer’s KV-cache, as its inversion directly yields the input sequence’s embeddings, allowing for precise mapping back to the vocabulary. In contrast, applying Eq. (4) to deeper layers recovers intermediate hidden states. These states represent semantically fused contextual information rather than discrete tokens. Reversing these deep representations into text is a non-trivial problem that necessitates training auxiliary inversion models [37], thereby increasing the adversary’s computational cost and reducing fidelity.
Consequently, while the Inversion Attack serves as a theoretical proof-of-concept for KV-cache leakage, its reliance on specific architectural properties (MHA) and layer positions (Layer 0) limits its utility against modern, deep LLMs. This motivates the need for the more robust attacks.